Event Type Date Time Logbook-Page

Failure 20-Jan-2002 20:55:00 hclog45730
Insert frozen into place due to missing he3 manometer plug.

Failure 21-Jan-2002 16:00:00 I-6
Target Mover adjusted.

Insert Change 21-Jan-2002 16:01:00 I-6
Insert C Top: None (Deut) Top: None (Deut)

Change 22-Jan-2002 22:00:00 I-8
Components installed for start of experiment.

Insert Change 23-Jan-2002 16:25:00 I-9
Insert Material Top: into (Deut) Top: (Deut)

TE Calibration 24-Jan-2002 00:21:51 I-11
Proton 10 samples 1.503K

TE Calibration 24-Jan-2002 01:54:52 I-11
Proton 20 samples 1.529K

TE Calibration 24-Jan-2002 04:47:51 I-13
Proton 20 samples 1.493K

Failure 24-Jan-2002 06:55:00 I-13
PDP Crash, perhaps lost TE data.

Anneal 26-Jan-2002 20:25:00 I-24
Empty Target 80K 15 minutes

Failure 27-Jan-2002 06:30:00 I-26
Run valve stuck.

Anneal 28-Jan-2002 09:17:00 I-31
Empty Target 80K 15 minutes

TE Calibration 29-Jan-2002 14:04:00 I-38
Proton 20 samples 1.547K

TE Calibration 29-Jan-2002 16:47:59 I-39
Proton 20 samples 1.551K

Anneal 29-Jan-2002 18:03:00 I-39
Empty Target 90K 15 minutes

Anneal 30-Jan-2002 08:46:00 I-41
Empty Target 90K 30 minutes

Failure 30-Jan-2002 11:30:00 I-43
Bypass valve stuck.

Anneal 30-Jan-2002 20:48:00 I-44
Empty Target 100K 15 minutes

Insert Change 31-Jan-2002 09:40:00 I-46
Insert B Top: E155_#11_scraps (Deut) Top: Gen01_Bottles_7_and_2 (Deut)

Change 01-Feb-2002 00:28:00 I-47
Lambda/2 cables changed.

Calibration 01-Feb-2002 08:57:00 I-48
Asymmetry study to get polarization

Anneal 01-Feb-2002 07:27:00 I-50
Empty Target 95K 60 minutes

TE Calibration 03-Feb-2002 11:50:00 I-54
Deuteron 8 samples 1.538K

TE Calibration 03-Feb-2002 13:05:00 I-55
Deuteron 10 samples 1.545K

TE Calibration 03-Feb-2002 14:29:00 I-56
Deuteron 10 samples 1.558K

Anneal 04-Feb-2002 08:40:00 I-60
Empty Target 85K 30 minutes

Anneal 04-Feb-2002 23:17:00 I-63
Empty Target 70K 10 minutes

Anneal 05-Feb-2002 11:56:00 I-66
Empty Target 75K 10 minutes

TE Calibration 05-Feb-2002 16:03:07 I-67
Deuteron 10 samples 1.540K

TE Calibration 05-Feb-2002 17:20:56 I-68
Deuteron 10 samples 1.531K

Calibration 05-Feb-2002 21:56:00 I-70
Decay study.

Failure 05-Feb-2002 02:00:00 I-71
Microwave counter failure.

Anneal 06-Feb-2002 12:21:00 I-72
Empty Target 90K 60 minutes

Anneal 07-Feb-2002 09:40:00 I-75
90K 60 minutes

Anneal 07-Feb-2002 18:46:00 I-78
Empty Target 90K 60 minutes

Failure 08-Feb-2002 08:15:00 I-80
Target motor control lost.

Anneal 08-Feb-2002 10:30:00 I-80
Empty Target 90K 60 minutes

Anneal 08-Feb-2002 21:55:00 I-83
100K 60 minutes

Failure 09-Feb-2002 02:30:00 I-85
Run valve stuck.

Anneal 09-Feb-2002 12:36:00 I-88
Carbon Target 95K 60 minutes

Anneal 09-Feb-2002 22:14:00 I-91
Empty Target 93K 60 minutes

Anneal 10-Feb-2002 08:11:00 I-93
Empty Target 95K 60 minutes

Anneal 10-Feb-2002 18:28:00 I-94
Empty Target 95K 60 minutes

Anneal 11-Feb-2002 07:27:00 I-96
Empty Target 95K 60 minutes

Anneal 11-Feb-2002 16:53:00 I-98
Empty Target 95K 60 minutes

Anneal 12-Feb-2002 00:17:00 I-99
Empty Target 100K 60 minutes

Change 12-Feb-2002 06:00:00 I-100
Magnet ramped down.

Calibration 12-Feb-2002 08:00:00 I-101
Table position encoder checked.

Change 12-Feb-2002 22:02:00 I-106
Magnet rotated.

Change 12-Feb-2002 22:02:00 I-106
Lots of changes made.

TE Calibration 13-Feb-2002 01:10:00 I-107
Deuteron 10 samples 1.544K

TE Calibration 13-Feb-2002 02:28:00 I-108
Deuteron 10 samples 1.544K

TE Calibration 13-Feb-2002 03:47:00 I-109
Deuteron 10 samples 1.563K

TE Calibration 13-Feb-2002 05:02:00 I-110
Deuteron 10 samples 1.567K

Insert Change 13-Feb-2002 12:14:00 I-111
Insert C Top: G04_G03_G08_G05 (Prot) Top: G10_&_E155_#9 (Prot)

TE Calibration 14-Feb-2002 17:16:00 I-113
Deuteron 20 samples 1.538K

TE Calibration 14-Feb-2002 19:37:00 I-114
Deuteron 20 samples 1.553K

Change 14-Feb-2002 20:38:00 I-115
Microwave frequency counter cable changed.

Failure 15-Feb-2002 07:05:00 I-116
Separator valve stuck.

Failure 16-Feb-2002 08:27:00 I-118
Heater wire possible too close to temp sensors, TCs not working.

Anneal 16-Feb-2002 08:30:00 I-118
85K 10 minutes

Anneal 16-Feb-2002 20:23:00 I-121
85K 15 minutes

Failure 17-Feb-2002 00:25:00 I-121
Run valve stuck open.

Failure 18-Feb-2002 05:30:00 I-124
Run valve stuck.

Anneal 18-Feb-2002 11:59:00 I-125
85K 15 minutes

Anneal 19-Feb-2002 09:50:00 I-129
85K 15 minutes

TE Calibration 20-Feb-2002 11:14:31 I-136
Proton 15 samples 1.529K

TE Calibration 20-Feb-2002 15:18:16 I-137
Proton 10 samples 1.522K

TE Calibration 20-Feb-2002 16:07:31 I-138
Proton 20 samples 1.532K

Insert Change 20-Feb-2002 21:00:00 I-138
Insert Top: (Deut) Top: (Deut)

TE Calibration 21-Feb-2002 01:55:37 I-139
Deuteron 13 samples 1.551K

TE Calibration 21-Feb-2002 03:34:21 I-139
Deuteron 15 samples 1.550K

TE Calibration 21-Feb-2002 05:54:45 I-140
Deuteron 9 samples 1.557K

Failure 21-Feb-2002 10:30:00 I-140
NMR bottom phase cable solder joint cracked.

Change 21-Feb-2002 10:40:00 I-140
Bottom NMR phase cable switched.

TE Calibration 21-Feb-2002 10:40:33 I-141
Deuteron 8 samples 1.549K

TE Calibration 21-Feb-2002 11:51:30 I-141
Deuteron 5 samples 1.537K

Anneal 22-Feb-2002 08:08:00 I-143
90K 60 minutes

Failure 22-Feb-2002 12:55:00 I-145
RB3 and RB2 tripped.

TE Calibration 22-Feb-2002 20:06:50 I-148
Deuteron 11 samples 1.525K

Anneal 24-Feb-2002 11:50:00 I-155
Empty Target 85K 25 minutes

Anneal 25-Feb-2002 09:39:00 I-158
Empty Target 75K 15 minutes

Failure 25-Feb-2002 10:43:00 I-158
Microwave counter not working.

Anneal 25-Feb-2002 20:42:00 I-159
80K 15 minutes

Anneal 26-Feb-2002 01:00:00 II-1
90K 15 minutes

TE Calibration 26-Feb-2002 11:47:00 II-2
Deuteron 10 samples 1.370K

TE Calibration 26-Feb-2002 13:03:00 II-3
Deuteron 7 samples 1.378K

TE Calibration 26-Feb-2002 14:09:00 II-3
Deuteron X samples XK

Anneal 26-Feb-2002 20:23:00 II-5
95K 20 minutes

Anneal 27-Feb-2002 11:01:00 II-7
95K 40 minutes

Anneal 28-Feb-2002 11:27:00 II-11
95K 60 minutes

Anneal 28-Feb-2002 15:58:00 II-12
100K 20 minutes

Failure 29-Feb-2002 01:40:00 II-13
Run valve problem.

Anneal 29-Feb-2002 08:58:00 II-14
95K 60 minutes

Anneal 02-Mar-2002 12:30:00 II-17
95K 45 minutes

Anneal 03-Mar-2002 08:15:00 II-20
95K 45 minutes

Anneal 03-Mar-2002 16:27:00 II-22
95K 60 minutes

TE Calibration 04-Mar-2002 02:33:35 II-23
Deuteron 10 samples 1.529K

TE Calibration 04-Mar-2002 03:44:00 II-24
Deuteron 10 samples 1.541K

TE Calibration 04-Mar-2002 06:34:00 II-26
Deuteron 28 samples XK

Change 04-Mar-2002 11:25:00 II-27
Experiment shut down.